# KAAE: Numerical Reasoning for Knowledge Graphs via Knowledge-aware Attributes Learning

### v1

@@ 4.3. Experimental Analysis
Figure 5. Proportional test of KAAE on the Credit dataset. A relation percentage of 100% means that we only consider relations, while 0% means that we only consider entities.
Proportional Analysis. Based on Eq. (3), we alter the proportion of weights between entities and relations in joint perception, exploring the improvement of numerical reasoning by the numerical information contained in entities and relations. Figure 5 illustrates the performance of KAAE when relations occupy different weights in joint perception, and the results show that the model achieves excellent performance when the proportion of relations is 40%-90%. There is an improvement of 46.1% in MRR and 60.2% in Hits@1 when the proportion of relations is 90% than when only entities are considered. However, the model performance deteriorates when we only consider relations and do not focus on entities. We can infer that both entities and relations contain information that is helpful for numerical reasoning. We need joint interaction of entity-relation attributes to accomplish numerical reasoning tasks efficiently.
Extreme value Analysis. To test whether our edge-range sampling approach will lead to a weaker perception of large numerical differences, we did an additional experiment to validate KAAE by selecting different proportions of positive/negative samples separately on Credit dataset. Table 5 demonstrates that KAAE can still maintain a good performance without utilizing all samples.
Table 5. Performance with different proportions of samples.
Proportion H@1 H@3 H@10 MR MRR
20% 0.739 0.805 0.882 7 0.796
40% 0.737 0.802 0.879 7 0.782
60% 0.746 0.821 0.885 7 0.789
80% 0.743 0.812 0.876 7 0.792
100% 0.745 0.818 0.888 7 0.794
Additionally, we added 10% extra extreme samples to the datasets. The results in Table 6 show that extreme values will not affect the performance of KAAE.
Table 6. Performance of KAAE with extra extreme values.
Dataset Extreme H@1 H@3 H@10 MR MRR
[…]
$\times$ 0.439 0.536 0.640 119 0.508
[…]
$\times$ 0.636 0.694 0.753 30 0.680
[…]
$\times$ 0.745 0.818 0.888 7 0.794
More Details for Experiments. For more details about the feasibility test, sensitivity experiment, loss functions, visualization results, and link evaluation, please refer to Appendix E.
5. Conclusion
In this paper, we introduce a new KGE model called KAAE. Specifically, to address the issue of semantic relevance, KAAE incorporates a joint interaction process that simultaneously captures the connections among entities, relations, and attributes within a unified framework. To improve semantic precision, we introduce a novel contrastive learning approach that enhances the diversity of training samples by optimizing ordinal samples based on cosine similarity and an entity-blending technique. Our KAAE model effectively enhances numerical reasoning in KGs. Extensive experiments on standard benchmarks such as US-Cities, Spotify, and Credit demonstrate KAAE’s effectiveness.
Appendix A Preliminaries
A.1. Knowledge Graph Embedding Methods
@@
Our model is based on the existing KGE method to model and evaluate the plausibility of the triples. Considering the continuity and scheduling problem of the numeric values, in this paper, we use TransE(Bordes et al., 2013) and Order-embedding(Vendrov et al., 2016). $\bm{e_{h}}$,$\bm{e_{h}}$, and $\bm{e_{r}}$ denote the embedding of head entities, tail entities, and relations respectively.The reasonableness score of the triple $(\bm{e_{h},e_{r},e_{t}})$ is obtained as follows.
@@
(17) $score(\bm{e_{h},e_{r},e_{t}})=\varepsilon-\|\bm{e_{h}+e_{r}-e_{t}}\|_{1/2},$
[…]
@@
Appendix B The Details of Proofs
B.1. The Loss of Contrastive Learning
We provide details of the derivations of KAAE in Eq. (12) and show that minimizing KAAE is approximately to minimize the combination of prediction and uniformity losses.
@@ B.1. The Loss of Contrastive Learning
(19) $\geq$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$
@@
(20) $\geq$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
B.2. Proof of Theorem
Minimizing KAAE’s objective in Equation (12) with an exponential moving average is equivalent to maximizing mutual information between representation $V$ and the one-hop pattern $Y$:
(22) $\begin{split}\mathcal{L}_{CL}\geq H(V|Y)-H(V)=-I(V;Y),\end{split}$
Proof. According to the derivations in Appendix B.1, we have:
@@ B.2. Proof of Theorem
[…]
@@
(24) $=$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\|\textbf{p}-\textbf{z}\|_{2}^{2}$
@@
where z $=\frac{1}{\mathcal{N}(v)}\sum\nolimits_{u\in\mathcal{N}(v)}$ u is the mean representations from one-hop neighborhoods of node $v$.Equation (23) holds because that gradient of u is being blocked in the target encoder with exponential moving average; thus it can be viewed as a constant that depends only on the target encoder but not the variable $f_{\theta}$. We can view the neighborhood representations $z$ as sampled from the conditional distribution given the pseudo label $Y$ which indicates the one-hop neighborhood pattern: $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$ because nodes belonging to the same semantic class tend to share similar one-hop neighborhood patterns. As demonstrated in [57], we can interpret the first term as a conditional cross-entropy between $V$ and another random variable $Z$ whose conditional distribution given the pseudo-label $Y$ indicates the one-hop neighborhood pattern: $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$:
[…]
(25) $=$ $H(V|Y)+\mathcal{D}_{KL}(V\|Z|Y)$
@@
where $H(\cdot)$ and $\mathcal{D}_{KL}(\cdot)$ denote denote the entropy and KL-divergence, respectively. Minimizing the first term in Equation (24) is approximately minimizing $H(V|Y)$. We then inspect the second term in Equation (24). As shown in [58, 57], the second term is close to the differential entropy estimator:
(26) $\frac{1}{|\mathcal{V}|}\frac{1}{|\mathcal{V}|}\sum\nolimits_{v\in\mathcal{V}}\sum\nolimits_{v_{-}\in\mathcal{V}}\|\textbf{v}-\textbf{v}_{-}\|_{2}^{2}\varpropto H(V),$
@@
As a result, minimizing the loss of KAAE can be seen as a proxy for maximizing the mutual information between the representations $V$ and the pseudo-labels $Y$ which indicates the neighborhood patterns of one hop nodes. Thus, the proof is completed.
Appendix C More Details of EXPERIMENTAL SETTINGS
The following subsections describe the data statistics, hyperparameters, experimental setup, and additional experimental results.
C.1. Dataset Statistics
@@
To evaluate the performance of KAAE in this paper, we used the US-Cities, Spotify, and Credit datasets, and detailed information about the datasets is described in Table 7 below.
Baseline. We classified various existing representation learning methods into six groups: Euclidean KGE including TransE(Bordes et al., 2013), ConvE(Dettmers et al., 2018), ComplEx (Trouillon et al., 2016), TuckER(Balazevic et al., 2019b), DistMult(Dong et al., 2014) and HAKE(Zhang et al., 2020). Hyperbolic KGE including MuRP(Balazevic et al., 2019a), ConE(Bai et al., 2021), AttH(Chami et al., 2020) and GIE(Cao et al., 2022b); GNNs-KGE including R-GCN(Schlichtkrull et al., 2018) and WGCN (Shang et al., 2019); Attributed KGE including KBLRN(García-Durán and Niepert, 2018), MT-KGNN (Tay et al., 2017) and LiteralE(Kristiadi et al., 2019); and Self-supervised model for graph including BiGI(Cao et al., 2021), SLiCE(Wang et al., 2021a), SimGCL(Yu et al., 2022), and RAKGE(Kim et al., 2023); Deterministic stand-alone value representation methods including DICE(Sundararaman et al., 2020), NEKG(Duan et al., 2021) and NRN(Bai et al., 2023).
  • Euclidean KGE. A representation learning method that models entity and relation embeddings in Euclidean space. TransE(Bordes et al., 2013) uses Euclidean distance as a distance measure. ConvE(Dettmers et al., 2018), ComplEx (Trouillon et al., 2016), TuckER(Balazevic et al., 2019b), and DistMult(Dong et al., 2014) use inner product as a distance measure. HAKE (Zhang et al., 2020) capture hierarchical relations. At the same time, the performance improvement strategy of LTE-KGE(Zhang et al., 2022) was implemented in all Euclidean KGE baselines above.
  • Hyperbolic KGE. Mapping KGs to specific hyperbolic spaces. This method operates well in circumstances where relations in KGs are hierarchical. MuRP(Balazevic et al., 2019a), ConE(Bai et al., 2021), and AttH(Chami et al., 2020) model hierarchical relations in latent space. Furthermore, GIE(Cao et al., 2022b) considers both hierarchical and ring-like structures.
  • GNNs-KGE. A method that makes use of GCN layers as entity encoders and predicts the plausibility score of each triple using the score functions of other KGE models. In this paper, we select two well-performing baselines: R-GCN(Schlichtkrull et al., 2018) and WGCN (Shang et al., 2019).
  • Attributed KGE. Utilizing side information to obtain more robust embeddings in sparsely observed KGs. KBLRN(García-Durán and Niepert, 2018), MT-KGNN (Tay et al., 2017) and LiteralE(Kristiadi et al., 2019) are well-performing KGE models based on numerical attributes.
  • Self-supervised model for graph. An approach that enriches node representation through self-supervised learning. BiGI(Cao et al., 2021) and SLiCE(Wang et al., 2021a) capture both global and local structures to enhance node representation. SimGCL(Yu et al., 2022) adjusts the uniformity of the learned representation by adding random noise vectors to node representations. RAKGE(Kim et al., 2023) captures the connection between relations and numerical attributes to improve node representation.
  • Deterministic value representation methods. To validate the effectiveness of our numerical embedding method, we explore the use of deterministic stand-alone value representations. DICE(Sundararaman et al., 2020)creates corpus-independent deterministic embeddings for numbers. NEKG(Duan et al., 2021) embeds the knowledge graph in a vector space using a graph embedding method for obtaining embeddings of the number entities. NRN(Bai et al., 2023) introduce new numerical variables and operations to describe queries involving numerical attribute values, alternatively encoding entities and numerical values into separate encoding structures.
Table 7. Data statistics
@@
C.2. Hyperparameters

### v3

@@ Introduction
Figure 4: Proportional test on Credit and Spotify. A relation percentage of 100% means that we only consider relations, while 0% means that we only consider entities.
Case Study. Furthermore, for easier visualization, we select part of the samples from Credit to show the attention scores between entities and attributes using heatmaps. As shown in Figure 5, the color blocks in each row of the first heat map are similar, indicating that RAKGE pays almost equal attention to each attribute in most samples, while the color blocks from NumCoKE are clearly differentiated, demonstrating that NumCoKE can better distinguish the importance of attributes in different relation contexts. More case studies can refer to Appendix E for more details.
Figure 5: Visualization of relevance scores of each numeric attribute of RAKGE and NumCoKE on Credit dataset.
Supplementary Experiments
We performed other insightful complementary experiments, including hyperparameter analysis, loss function experiment, evaluation on FB15k-237, etc. (see Appendix E).
Conclusion
In this paper, we introduce a novel KGE model called NumCoKE. Specifically, we incorporates a Mixture-of-Experts-Knowledge-Aware Encoder for elaborate semantic modeling, capturing the connections among entities, relations, and attributes within a unified framework. To distinguish similar semantics, we introduce a novel ordinal knowledge contrastive learning strategy that that generates high-quality ordinal samples from original data, capturing fine-grained semantic nuances over close numerical values. Extensive experiments on three standard benchmarks demonstrate NumCoKE’s effectiveness for numerical reasoning on KGs.
Acknowledgments
This work was supported by the National Key R&D Program of China under Grant 2022YFB3903904, the Key R&D Program of Jiangxi Province under Grant 20232BBGW001.
Catalog Content
A Knowledge Graph Embedding Methods
B The Details of Proofs
B.1 The loss of Contrastive Learning
B.2 Proof of Theorem
C Experiment Setup
C.1 Dataset Statistics
C.2 Baselines
C.3 Hyperparameter Settings
D Learning Procedure (Algorithm)
E Supplementary Experiments
E.1 Single Module Ablation Study
E.2 Hyperparameter Sensitivity Experiment
E.3 Loss Function Experiment
E.4 Feasibility Test
E.5 Extreme value Analysis
E.6 Backbone Model Analysis
E.7 Visualization
E.8 Evaluation on FB15k-237 Dataset
E.9 Breakdown Analysis
E.10 Case Study
Table 3: Appendix summary table, including indexes and content titles for each chapter.
Appendix A A Knowledge Graph Embedding Methods
@@
Our model is based on the existing KGE method to model and evaluate the plausibility of the triples. Considering the continuity and scheduling problem of the numeric values, in this paper, we use TransE (DBLP:conf/nips/BordesUGWY13) and Order-embedding (DBLP:journals/corr/VendrovKFU15). $\bm{e_{h}}$$\bm{e_{h}}$, and $\bm{e_{r}}$ denote the embedding of head entities, tail entities, and relations respectively. The reasonableness score of the triple $(\bm{e_{h},e_{r},e_{t}})$ is obtained as follows.
@@
$score(\bm{e_{h},e_{r},e_{t}})=\varepsilon-\|\bm{e_{h}+e_{r}-e_{t}}\|_{1/2},$ (18)
where $\varepsilon$ is a hyperparameter and $\|\cdot\|$ can be L1 or L2 distance.
@@
Appendix B B The Details of Proofs
B.1 The Loss of Contrastive Learning
We provide details of the derivations in Eq. (13) and show that minimizing NumCoKE is approximately to minimize the combination of prediction and uniformity losses.
@@ B.1 The Loss of Contrastive Learning
$\geq$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$ (20)
@@
$\geq$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$ (21)
@@
$\overset{c}{=}$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$
@@
$\overset{c}{=}$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\textbf{p}^{\top}\textbf{u}$
@@
$\overset{c}{=}$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}\|\textbf{p}-\textbf{u}\|_{2}^{2}$ (22)
@@
where the symbol $\overset{c}{=}$ indicates equality up to a multiplicative and/or additive constant. Here, we utilize Jensen’s inequality in Equation (21). Equation (22) holds because p and u are both normalized. Given the above, we can conclude that minimizing NumCoKE objective is approximately to minimize combination of the prediction and uniformity losses.
B.2 Proof of Theorem
Minimizing NumCoKE’s objective in Equation (13) with an exponential moving average is equivalent to maximizing mutual information between representation $V$ and the one-hop pattern $Y$:
$\begin{split}\mathcal{L}_{CL}\geq H(V|Y)-H(V)=-I(V;Y),\end{split}$ (23)
Proof. According to the derivations in Appendix B, we have:
@@ B.2 Proof of Theorem
$\overset{c}{=}$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\|\textbf{P}-\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}\textbf{u}\|_{2}^{2}$ (24)
@@
$=$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\|\textbf{p}-\textbf{z}\|_{2}^{2}$ (25)
@@
where z $\frac{1}{\mathcal{N}(v)}\sum\nolimits_{u\in\mathcal{N}(v)}$ u is the mean representations from one-hop neighborhoods of node $v$.Equation (24) holds because that gradient of u is being blocked in the target encoder with exponential moving average; thus it can be viewed as a constant that depends only on the target encoder but not the variable $f_{\theta}$. We can view the neighborhood representations $z$ as sampled from the conditional distribution given the pseudo label $Y$ which indicates the one-hop neighborhood pattern: $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$ because nodes belonging to the same semantic class tend to share similar one-hop neighborhood patterns. As demonstrated in (DBLP:journals/jmlr/WangS11), we can interpret the first term as a conditional cross-entropy between $V$ and another random variable $Z$ whose conditional distribution given the pseudo-label $Y$ indicates the one-hop neighborhood pattern: $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$:
$\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\|\textbf{p}-\textbf{z}\|_{2}^{2}\overset{c}{=}$ $H(V;Z|Y)$
$=$ $H(V|Y)+\mathcal{D}_{KL}(V\|Z|Y)$ (26)
@@
where $H(\cdot)$ and $\mathcal{D}_{KL}(\cdot)$ denote denote the entropy and KL-divergence, respectively. Minimizing the first term in Equation (25) is approximately minimizing $H(V|Y)$. We then inspect the second term in Equation (25). As shown in (DBLP:conf/eccv/BoudiafRZGPPA20; DBLP:journals/jmlr/WangS11), the second term is close to the differential entropy estimator:
$\frac{1}{|\mathcal{V}|}\frac{1}{|\mathcal{V}|}\sum\nolimits_{v\in\mathcal{V}}\sum\nolimits_{v_{-}\in\mathcal{V}}\|\textbf{v}-\textbf{v}_{-}\|_{2}^{2}\varpropto H(V),$ (27)
@@
As a result, minimizing the loss of NumCoKE can be seen as a proxy for maximizing the mutual information between the representations $V$ and the pseudo-labels $Y$ which indicates the neighborhood patterns of one hop nodes. Thus, the proof is completed.
Appendix C C Experiment Setup
The following section describes the experiment setup, including dataset statistics, the selection of baseline methods, and hyperparameter settings.
C.1 Dataset Statistics
@@
To evaluate the performance of NumCoKE, we use the US-Cities, Spotify, and Credit datasets, and detailed information about the datasets is described in Table 4 below.
US-Citiesis a large-scale city dataset containing basic information about cities in the United States, including 65 attribute fields (e.g., household_income, commute_time, etc). Spotify is a knowledge graph provided by Spotify for developers.Credit is a knowledge graph constructed from default events in Taiwan (DBLP:journals/eswa/YehL09a). Each credit card holder corresponds to an entity with numerical information about the credit card (e.g., total_delay, total_bill) as the numerical attribute. Additionally, all the three datasets are modified to simulate numerical reasoning tasks under real conditions. Close to 20% of the numerical values are masked to zero (missing values) while using as few triples as possible for the training set. The remaining triples are used for evaluation.
@@
Table 4: Data statistics
C.2 Baselines
We classified various existing representation learning methods into six groups: Euclidean KGE including TransE (DBLP:conf/nips/BordesUGWY13), ConvE (DBLP:conf/aaai/DettmersMS018), TuckER (DBLP:conf/emnlp/BalazevicAH19), HAKE (DBLP:conf/aaai/ZhangCZW20). DaBR (DBLP:conf/coling/WangLBG25), KGDM (DBLP:conf/aaai/LongZLWLW24), and RuIE (DBLP:conf/coling/LiaoDH025). Hyperbolic KGE including MuRP (DBLP:conf/nips/BalazevicAH19), ConE (DBLP:conf/nips/BaiYRL21), AttH (DBLP:conf/acl/ChamiWJSRR20), GIE (DBLP:conf/aaai/CaoX0CH22) and LorentzKG (DBLP:conf/acl/Fan0CCDY24); GNNs-KGE including R-GCN (DBLP:conf/esws/SchlichtkrullKB18) and WGCN (DBLP:conf/aaai/ShangTHBHZ19); Attributed KGE including KBLRN (DBLP:conf/uai/Garcia-DuranN18), MT-KGNN (DBLP:conf/cikm/TayTPH17) and LiteralE (DBLP:conf/semweb/KristiadiKL0F19); and Self-supervised model for graph including BiGI (DBLP:conf/wsdm/CaoLGLL021), SLiCE (DBLP:conf/www/WangAHCR21), SimGCL (DBLP:conf/sigir/YuY00CN22), and RAKGE (DBLP:conf/kdd/KimKKPJP23); Deterministic stand-alone value representation methods including DICE (DBLP:conf/emnlp/SundararamanSSW20), NEKG (DBLP:conf/emnlp/DuanYT21) and NRN (DBLP:conf/kdd/BaiLLYYS23). We reproduced all baselines with official settings and tuning to ensure a fair comparison.
  • Euclidean KGE. A representation learning method that models entity and relation embeddings in Euclidean space. TransE (DBLP:conf/nips/BordesUGWY13) uses Euclidean distance as a distance measure. ConvE (DBLP:conf/aaai/DettmersMS018), TuckER (DBLP:conf/emnlp/BalazevicAH19) use inner product as a distance measure. HAKE (DBLP:conf/aaai/ZhangCZW20) capture hierarchical relations. DaBR (DBLP:conf/coling/WangLBG25) utilize distance-adaptive translations to learn geometric distance between entities, KGDM (DBLP:conf/aaai/LongZLWLW24) estimates the probabilistic distribution of target entities in prediction through diffusion models. RuIE (DBLP:conf/coling/LiaoDH025) leverage logical rules to enhance reasoning. At the same time, the performance improvement strategy of LTE-KGE (DBLP:conf/www/Zhang0YW22) was implemented in all Euclidean KGE baselines above.
  • Hyperbolic KGE. Mapping KGs to specific hyperbolic spaces. This method operates well in circumstances where relations in KGs are hierarchical. MuRP (DBLP:conf/nips/BalazevicAH19), ConE (DBLP:conf/nips/BaiYRL21), and AttH (DBLP:conf/acl/ChamiWJSRR20) model hierarchical relations in latent space. Furthermore, GIE (DBLP:conf/aaai/CaoX0CH22) considers both hierarchical and ring-like structures. LorentzKG (DBLP:conf/acl/Fan0CCDY24) embed KGs into hyperbolic spaces to model hierarchical relations.
  • GNNs-KGE. Methods make use of GCN layers as entity encoders and predict the plausibility score of triples using score functions of other KGE models. We select two well-performing baselines: R-GCN (DBLP:conf/esws/SchlichtkrullKB18) and WGCN (DBLP:conf/aaai/ShangTHBHZ19).
  • Attributed KGE. Utilizing side information to obtain more robust embeddings in sparsely observed KGs. KBLRN (DBLP:conf/uai/Garcia-DuranN18), MT-KGNN (DBLP:conf/cikm/TayTPH17) and LiteralE (DBLP:conf/semweb/KristiadiKL0F19) are well-performing models based on numerical attributes.
  • Self-supervised model for graph. Methods that enriches node representation through self-supervised learning. BiGI (DBLP:conf/wsdm/CaoLGLL021) and SLiCE (DBLP:conf/www/WangAHCR21) capture both global and local structures to enhance node representation. SimGCL (DBLP:conf/sigir/YuY00CN22) adjusts the uniformity of the node representation by adding random noise vectors. RAKGE (DBLP:conf/kdd/KimKKPJP23) captures the connection between relations and numerical attributes to improve node representation.
  • Deterministic value representation methods. To validate our numerical embedding method, we explore the use of deterministic stand-alone value representations. DICE (DBLP:conf/emnlp/SundararamanSSW20) creates corpus-independent deterministic embeddings for numbers. NEKG (DBLP:conf/emnlp/DuanYT21) embeds the knowledge graph in a vector space using a graph embedding method for obtaining embeddings of the number entities. NRN (DBLP:conf/kdd/BaiLLYYS23) introduce new numerical variables and operations to describe queries involving numerical attribute values, alternatively encoding entities and numerical values into separate encoding structures.
C.3 Hyperparameter Settings